A. Solve the following:

1. If the price increases from 30 to 40 and the quantity demanded falls from 50 to 30, what is the elasticity of demand?

2. If the price decreases from 20 to 10 and quantity demanded rises from 10 to 50 units, what is the elasticity of demand?


1
Expert's answer
2021-10-14T11:37:26-0400

1. "Ed = \frac{30-50} {40-30} ×\frac{40+30} {30+50} = -1.75."

So, the demand is elastic.

2. "Ed = \frac{50-10} {10-20} ×\frac{10+20} {50+10} = -2."

So, the demand is elastic.